# Conceptualization, Formal analysis, Investigation, Methodology, Software, Data curation, Validation, Visualization, Writing –original draft, Writing – review & editing

### v3

@@ 2 Related Works
Extrinsic calibration methods for LiDAR–camera systems can generally be divided into two categories: target-based and targetless approaches. Target-based calibration relies on specially designed calibration targets and is thus commonly associated with manual, offline procedures, although some automated methods do exist. In contrast, targetless methods extract features directly from natural scenes, making them well-suited for automatic, online calibration.
2.1 Target-based Calibration
Target-based calibration methods rely on the use of custom-designed calibration targets to extract corresponding features from the data of different sensors, enabling the calculation of extrinsic parameters. Beltrán et al. (Beltrán et al., 2022) propose a method for calibrating extrinsic parameters between LiDARs and cameras (monocular or stereo) of different modalities. The approach involves two stages: extracting reference points from a custom-designed calibration target and determining the optimal rigid transformation by registering these point sets. The calibration target, fabricated using a CNC machine, features four circular holes to utilize geometric discontinuities in LiDAR and stereo point clouds, along with four ArUco markers placed at the corners for inferring 3D information from monocular images. Fu et al. (Fu et al., 2019) propose a LiDAR–camera calibration method using multiple chessboards placed in various poses. A stationary LiDAR scan captures the chessboards, and the camera is moved to obtain sequential images for reconstructing a 3D visual point cloud. The extrinsic parameters are then calculated by aligning the 3D visual points with the LiDAR scan using a tightly coupled graph optimization method. Zhang et al. (Zhang et al., 2023a) introduce the $L^{2}V^{2}T^{2}Calib$ method, which uses a four-circular-holes board as a calibration target. The centers of the circles serve as common features, detectable by sparse and dense LiDAR, repetitive and non-repetitive scanning LiDAR, as well as visual and thermal cameras. The method employs template matching for automatic target detection and calculates the calibration matrix by minimizing the 2D reprojection error. Koo et al. (Koo et al., 2022) propose a method to formulate analytic plane covariances from a 2D checkerboard planar target image, quantifying the weighting of objective functions and enhancing camera–LiDAR extrinsic calibration performance. The method uses a typical objective function based on plane feature correspondences between the camera and LiDAR. For each checkerboard image, the 3D transformation between the camera and checkerboard coordinate frames is computed using the checker square length and the PnP algorithm. Li et al. (Li et al., 2022) address the challenge of establishing common feature correspondences between sparse LiDAR point clouds and monocular images by designing a novel calibration board featuring checkerboard grids and circular holes. The extrinsic parameters are automatically determined by matching the centers of the circular holes extracted from both images and heterogeneous LiDAR scans. Huang et al. (Huang et al., 2024) propose a novel calibration method using a uniquely designed acrylic checkerboard that allows LiDAR beams to pass through the white grids and reflect off the black grids. This design reduces the size of the calibration board, enabling calibration at close distances while ensuring sufficient corresponding features can be extracted from both LiDAR scans and images to achieve accurate calibration results. Itami et al. (Itami and Yamazaki, 2020) propose an improved calibration method for a 2D LiDAR and camera system, utilizing a simple checkerboard target and a vertically rotating platform. This approach addresses a common drawback in LiDAR calibration with checkerboards—the sensitivity to checkerboard poses—by enabling more consistent and robust calibration results. Toth et al. (Tóth et al., 2020) introduce an automatic calibration method using spheres for LiDAR and camera calibration. Sphere centers are independently determined from LiDAR point clouds using RANSAC and LSQ regression and from camera images via edge detection and ellipse fitting. The extrinsic parameters of the LiDAR–camera pair are then determined through center point registration. Verma et al. (Verma et al., 2019) propose an automated method for extrinsic calibration between a camera and LiDAR using a checkerboard as a reference. The calibration board’s center point and normal vector are automatically extracted from the LiDAR point cloud by leveraging its geometric properties. Corresponding features in the camera image are derived from the camera’s extrinsic matrix. Once feature correspondences are established, a Genetic Algorithm is employed to achieve a globally optimal calibration result.
Despite their high precision, target-based methods have notable limitations. They require meticulous target preparation, often involving custom-designed and precisely fabricated targets, as well as specific setups in controlled environments. These constraints reduce their flexibility and scalability, particularly in dynamic or large-scale scenarios. Moreover, these methods generally cannot handle runtime decalibrations, a frequent occurrence in real-world applications, potentially compromising the accuracy and reliability of the calibration results.
2.2 Targetless Calibration
Targetless methods eliminate the need for manually placed dedicated calibration targets by detecting and matching naturally occurring features observable by both sensors to establish correspondences, making them more suitable for fully automatic online calibration. Depending on the type of features (or discriminative information) extracted from the scene, these approaches can be broadly categorized into motion-based, edge alignment-based, mutual information-based, and deep learning–based methods.
2.2.1 Motion-based Calibration
Petek et al. (Petek et al., 2024) introduce MDPCalib, an automated camera–LiDAR extrinsic calibration approach that requires no dedicated targets. Their method employs visual and LiDAR odometry to generate two sensor motion paths, which are aligned via a non-linear optimization for coarse registration. This coarse alignment then initializes a learning-based 2D–3D point correspondence algorithm, yielding dense matches between image and point cloud spaces. Finally, camera–LiDAR calibration is formulated as an optimization problem that minimizes costs derived from sensor motion and point correspondences, resulting in accurate extrinsic parameters. Park et al. (Park et al., 2020) propose a LiDAR-camera calibration method that leverages each sensor’s motion, which is estimated via LiDAR odometry and visual odometry. From the resulting trajectories, sets of relative transformations are extracted. By analyzing these pairs of relative transformations in both the LiDAR and camera coordinate frames, the method derives a closed-form solution for the LiDAR–camera extrinsic calibration. Yoon et al. (Yoon et al., 2021) present a targetless approach for calibrating multiple camera-LiDAR systems by leveraging object pose estimation. First, an initial calibration is generated by registering the LiDAR point cloud with an up-to-scale Structure-from-Motion (SfM) point cloud, linking corresponding instance segments in the image and point cloud domains. Based on these correspondences, a color appearance model is constructed for each object. Subsequently, iterative region-based object pose estimation is performed using the initial pose, mesh, and color appearance models, thereby refining the extrinsic parameters for multiple sensors. Some motion-based (or pose-based) methods rely on hand-eye calibration. For example, Ou et al. (Ou et al., 2023) propose a targetless LiDAR–camera calibration method leveraging cross-modality structure consistency to address the degeneration issue of hand-eye calibration when sensor motions lack sufficient rotation. Their approach uses visual and LiDAR SLAM to estimate camera and LiDAR poses from the collected data. Hand-eye calibration is then applied to compute an initial extrinsic matrix, which is further refined through a global optimization process to achieve the optimal extrinsic parameters.
These methods, while effective in leveraging sensor motion or pose data for extrinsic calibration, have notable limitations. They rely heavily on accurate motion estimation, which can be challenging in scenarios with noisy data or insufficient sensor motion, such as limited rotation. Additionally, these methods often require complex preprocessing steps, such as odometry or SLAM-based pose estimation, which may introduce errors or computational overhead. Furthermore, some approaches that rely on hand-eye calibration are prone to degeneration under specific motion patterns, such as linear trajectories with minimal rotational movement, further limiting their applicability in diverse real-world scenarios.
2.2.2 Edge Alignment-based Calibration
Zhang et al. (Zhang and Rajan, 2022) propose Multi-FEAT (Multi-Feature Edge AlignmenT), a targetless approach for camera-LiDAR extrinsic calibration. The method transforms the 2D (camera)–3D (LiDAR) calibration problem into a 2D–2D calibration problem using a cylindrical projection model. By leveraging various LiDAR features, it reconstructs edges from the sparse LiDAR point cloud more effectively. A cost function is designed to align the edge intensities of camera image edges with the edge probabilities derived from the multi-feature LiDAR point cloud. The unknown extrinsic parameters are then estimated using a gradient ascent optimization method. Li et al. (Li et al., 2024) introduce EdgeCalib, an edge-based method for automatic targetless calibration of LiDARs and cameras in real-world scenarios. The approach utilizes edge features to establish reliable correspondences between images and point clouds. The Segment Anything Model is employed to extract stable and robust image edge features, while a multi-frame weighting strategy filters features and reduces environmental dependency. Accurate extrinsic parameters are then estimated by enforcing constraints on the edge correspondences. Yuan et al. (Yuan et al., 2021) propose an automatic extrinsic calibration method designed for targetless environments. The system extracts natural edge features from both images and point clouds and minimizes the reprojection error to achieve calibration. The method includes an analysis of the constraints imposed by edge features and evaluates the impact of edge distribution on calibration accuracy. To enhance performance, they introduce an efficient LiDAR edge extraction technique based on point cloud voxel cutting and plane fitting.
Edge alignment-based calibration methods face inherent challenges due to the difficulty of matching edges between LiDAR and camera data. Object boundaries are often arbitrary, making it challenging to define consistent and reliable edge features. Furthermore, the differences in sensing modalities result in significant variations in how edges are represented, as LiDAR captures sparse geometric structures while cameras capture dense visual information. These discrepancies can lead to mismatched features, reducing the accuracy and robustness of the calibration process.
2.2.3 Mutual Information–based Calibration
Pandey et al. (Pandey et al., 2012) address the automatic, targetless extrinsic calibration of a 3D LiDAR and camera system using a mutual information (MI) framework. The algorithm registers intensity information from the camera with reflectivity information from the LiDAR. Calibration parameters are estimated by maximizing the mutual information between the sensor-measured surface intensities, enabling accurate cross-modal alignment. Taylor et al. (Taylor and Nieto, 2013) propose an automatic calibration method for camera–LiDAR systems using normalized mutual information. The approach compares camera images with LiDAR scans of the same area. A camera model incorporating orientation, location, and focal length is used to generate a 2D LiDAR image, where pixel intensities represent features of the LiDAR scan. Particle swarm optimization is then applied to maximize the normalized mutual information and determine the optimal calibration parameters.
Mutual information–based calibration methods have significant limitations due to their reliance on reflectance intensities. LiDAR reflectance can vary widely depending on the material properties and surface characteristics, while camera pixel intensities are highly sensitive to lighting conditions. These variations can lead to unreliable correspondences, limiting the effectiveness and accuracy of such calibration approaches.
2.2.4 Deep Learning-based Calibration
Schneider et al. (Schneider et al., 2017) introduce RegNet, the first convolutional neural network (CNN) designed to directly regress the extrinsic calibration between sensors of different modalities. The problem is framed as determining the decalibration between an initial calibration matrix and a ground truth calibration matrix. To establish correspondences, LiDAR points are projected onto the camera frame using the initial calibration matrix, creating a projected LiDAR depth image. RegNet infers the correspondence between these projected depth measurements and the RGB image, ultimately regressing the extrinsic calibration parameters. Subsequent deep learning-based methods mostly follow the RegNet paradigm. For example, Lv et al. (Lv et al., 2021) propose LCCNet (LiDAR-Camera Self-Calibration Network), an online calibration framework. LCCNet employs a cost volume layer to capture the correlation between RGB image features and the depth image projected from point clouds. The network takes as input an RGB image from a calibrated camera and a projected sparse depth image from a mis-calibrated LiDAR. It outputs a 6-DoF rigid-body transformation, representing the deviation between the initial extrinsic matrix and the ground truth extrinsic matrix. Iyer et al. (Iyer et al., 2018) introduce CalibNet, a geometrically supervised deep network for real-time estimation of the 6-DoF rigid body transformation between a 3D LiDAR and a 2D camera. The network features a novel architecture based on 3D Spatial Transformers, which solves the calibration problem by maximizing geometric and photometric consistency. It takes as input an RGB image and a sparse depth map generated from a mis-calibrated LiDAR point cloud. The training process incorporates geometric supervision by minimizing dense photometric error and dense point cloud distance error, enabling the network to accurately regress the extrinsic calibration parameters. Shi et al. (Shi et al., 2020) propose CalibRCNN (Calibration Recurrent Convolutional Neural Network) to infer a 6-DoF rigid body transformation between 3D LiDAR and 2D camera. It uses an LSTM network to extract temporal features from consecutive frames of 3D point clouds and RGB images and refines calibration accuracy with geometric and photometric losses from interframe constraints. By leveraging the correspondence between projected LiDAR depth images and RGB images, CalibRCNN learns the underlying 2D–3D geometry for improved calibration precision. Xiao et al. (Xiao et al., 2024) propose CalibFormer, an end-to-end network for automatic LiDAR–camera calibration. The approach projects the LiDAR point cloud onto the image plane using the initial extrinsic parameter and camera matrix to generate a miscalibrated LiDAR image. The network takes both camera images and LiDAR images as inputs, applying a multi-head correlation module to compute correspondences between misaligned features across different dimensions. A transformer architecture is then used to extract and enhance high-contribution correlation features. Finally, the network regresses the deviations of the calibration parameters. Some deep learning methods diverge from the RegNet paradigm and instead utilize semantic segmentation models for LiDAR–camera calibration. For example, Luo et al. (Luo et al., 2024) propose a calibration method leveraging the Segment Anything Model (SAM) without additional training. SAM is used to automatically generate segmentation masks for input images. For the point cloud, normal estimation, clustering, and intensity normalization are applied to assign attributes to each point, including intensity, normal vector, and segmentation class. The extrinsic parameters are then optimized by maximizing the consistency score of the point attributes that correspond to each mask.
Current deep learning–based calibration methods, while innovative, have notable limitations. Approaches following the RegNet paradigm require an initial calibration to project LiDAR data into the image plane, creating a "projected LiDAR depth map." This reliance on initial calibration, often obtained through manual intervention or empirical estimates, reduces the method’s generalizability. Additionally, the use of projected depth maps overemphasizes image data, as the sparse nature of LiDAR point clouds results in extracted features being dominated by image characteristics, sidelining valuable LiDAR-specific information. Furthermore, directly regressing calibration matrices presents a computational challenge, akin to solving an unconstrained task within constrained optimization, which challenges the capability of current neural networks. These models also face significant computational overhead, making them unsuitable for real-time applications. Methods based on semantic segmentation models also have drawbacks, as they often focus solely on image-based segmentation. This approach is computationally expensive and fails to provide a significant advantage over object detection methods for calibration tasks, further questioning their practicality.
Moreover, most existing methods fail to fully exploit the advances in object detection and feature extraction developed for LiDAR and camera data processing. In contrast to existing methods, we propose a fully automatic, targetless, and online calibration framework that directly processes raw LiDAR point clouds and camera images, eliminating the need for initial calibration, projected depth maps, or heuristic preprocessing. Our approach builds on proven object detection algorithms to effectively extract objects from both sensor modalities and introduces a Common Feature Discriminator to establish correspondences by learning and matching features such as relative positions, appearance embeddings, and classification information across the two modalities. To ensure high accuracy, we introduce a two-stage refinement process: iterative refinement continuously refines the calibration by leveraging newly established correspondences, while an attention-based refinement employs a Vision Transformer and cross-attention mechanisms to address complex challenges such as non-planar surfaces, depth variations, and sensor misalignments to further refine the calibration. This method enables the direct matching of corresponding points, facilitating a straightforward, one-shot, and end-to-end calibration process between the LiDAR and camera, with improved adaptability to real-world scenarios.
@@ 2.2.4 Deep Learning-based Calibration
3 Proposed Method
3.1 Problem Formulation
Extrinsic calibration between sensors aims to unify detections from two different sensors into the same frame of reference or coordinate system, enabling the fusion of their respective detection information. For instance, a single target appearing in both the LiDAR and camera views can provide complementary details—such as color from the camera and 3D shape from the LiDAR. Additionally, LiDAR can detect objects at greater distances that a camera might miss, compensating for the camera’s detection limitations. Consequently, calibration allows for a more comprehensive and robust perception by combining the strengths of both sensors.
LiDAR–camera extrinsic calibration is typically accomplished by solving for a transformation matrix that associates a point in the image pixel coordinate system (PCS) with its corresponding point in the LiDAR coordinate system (LCS). Since points in the LCS are 3D, while those in the PCS are 2D, most existing calibration methods rely on 3D-to-2D perspective projection. However, this approach has notable drawbacks. First, it requires the camera’s intrinsic matrix, adding the burden of intrinsic camera calibration, which is often performed manually, thus hindering fully automatic extrinsic calibration. Second, estimating the 3D-to-2D transformation matrix is computationally more complex and prone to instability. More importantly, for most practical applications, 3D-to-2D perspective calibration is unnecessary for achieving effective LiDAR–camera data fusion. A simpler 2D-to-2D projective calibration, where the 2D LiDAR plane is obtained by removing the Z-axis, is sufficient.
This simplification is justified for several reasons. The primary goal of calibration is to enable data fusion between the two sensors, such as associating 3D LiDAR point cloud clusters with image pixel regions for the same object. Achieving this does not require projecting the 3D LiDAR point cloud onto the image plane using a 3D-to-2D calibration matrix. Most existing methods adopt the 3D-to-2D approach as it draws from camera calibration practices that focus on 3D reconstruction. However, LiDAR–camera calibration is fundamentally different, as its focus is on data fusion, not reconstruction. By using 2D-to-2D projective calibration, where 2D LiDAR points are mapped to the image plane, corresponding 3D LiDAR points can still be retrieved without requiring a 3D-to-2D perspective transformation. Additionally, when projecting 3D LiDAR points onto the image plane, the LiDAR data effectively becomes 2D, resulting in the loss of LiDAR’s inherent 3D detection capabilities. Thus, 3D-to-2D calibration does not provide more useful information than 2D-to-2D projective calibration. Notably, while many existing methods emphasize projecting 3D LiDAR points onto the image plane, this should only serve as a visualization tool to intuitively present calibration performance, not as the calibration objective itself. The true goal of calibration should be the seamless and accurate fusion of sensor data.
Thus, we propose using planar projective transformation to achieve 2D-to-2D calibration between the 2D LiDAR plane and the camera image plane, as illustrated in Fig. 3. A planar projective transformation, or Homography, is an invertible linear transformation represented by a non-singular matrix $\mathbf{H}\in\mathbb{R}^{3\times 3}$ (Dubrofsky, 2009). This transformation allows us to project a point in the LCS directly onto the camera image plane without requiring the camera intrinsic matrix. The relationship is expressed as:
@@ 3.1 Problem Formulation
where $\hat{P_{l}}=(u_{l},v_{l})$ is the projection of a point $P_{l}=(x,y)$ in the LCS onto the camera image plane PCS. Notably, the objects or points on the 2D LiDAR plane and those on the camera image plane are derived from objects or points lying on a common plane (e.g., the ground plane), as shown in Fig. 3. This alignment justifies the use of 2D Homography for LiDAR–camera calibration, as it can be considered a planar homography induced by the common plane (Szeliski, 2022). To solve for the Homography matrix, a set of $N$ points in the LCS and their corresponding points in the PCS is required. While 4 points are theoretically sufficient to compute the matrix, more points are typically used to optimize the solution based on a specific cost function (Dubrofsky, 2009). The most widely used cost function is the minimization of the geometric reprojection error (Hartley and Zisserman, 2003; Dubrofsky, 2009), which evaluates the alignment between $N$ projected 2D LiDAR pixel points ($\hat{P_{l}}=(\hat{u},\hat{v})$) and their corresponding 2D image pixel points ($P_{p}=(u,v)$). Minimizing this reprojection error ensures an accurate Homography matrix for precise calibration. Specifically, the reprojection error can be measured as either the average Euclidean distance (L2 norm mean) or the root mean square error.
1. Average Euclidean Distance (AED): It computes the mean of the Euclidean distances between the projected points and their corresponding ground truth points. It provides a straightforward and intuitive measure of overall alignment, offering insight into the average deviation between matched points:
@@
$\mathcal{E}_{\text{AED}}=\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{v}^{i})^{2}}.$ (2)
2. Root Mean Square Error (RMSE): Alternatively, the reprojection error can be expressed as the root mean square error, which is more sensitive to larger deviations. By squaring the Euclidean distances before averaging, this measure amplifies the influence of outliers, ensuring that significant misalignments are effectively penalized:
@@
$\mathcal{E}_{\text{RMSE}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}^{2}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left[(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{u}^{i})^{2}\right]}.$ (3)
3.2 Method Overview
We aim to develop a framework for LiDAR–camera online automatic targetless calibration, designed to reduce human intervention, streamline sensor integration, and ensure high precision in LiDAR–camera fusion applications. Our method comprises the following stages, as shown in Fig. 1:
  1. 1.
    Stage 1: In this stage, the established LiDAR and camera detectors are utilized to detect and extract objects from their respective sensor data. Each detected object is extracted, including its center position (the bounding box center for the camera detection and the cluster center for the LiDAR point cloud). These objects are then used to train a Common Feature Discriminator, which determines whether an image object and a LiDAR object correspond to the same entity. To achieve this, the discriminator learns and compares three distinct features: Relative Positions, Appearance Embeddings, and Classification Information. These features are concatenated and passed through a feed-forward neural network (FFN) classifier, which outputs a decision on whether the objects from the two sensors are the same or different.
  2. 2.
    Stage 2: This stage involves solving the calibration matrix using the identified object pairs. A homography transformation is applied to generate a coarse initial calibration matrix ($H$), which establishes a preliminary correspondence between objects detected by the LiDAR and camera.
  3. 3.
    Stage 3: In this stage, an iterative refinement-based fine calibration is performed to refine the calibration matrix, considering that the Common Feature Discriminator may not precisely match all corresponding objects. The initial calibration matrix from the previous stage is used to project LiDAR data onto the camera plane, enabling the selection and construction of additional point correspondences based on distance criteria. These newly established correspondences are then used to achieve more precise calibration through iterative refinement.
  4. 4.
    Stage 4: In the final stage, attention-based refinement employs a Vision Transformer to extract global distortion features from images, effectively addressing the limitations of homography calibration caused by non-planar surfaces and depth variations. This also compensates for errors introduced by the absence of camera intrinsic matrix-based rectification. Furthermore, it integrates a cross-attention network to compute weighted interactions between image pixels (as queries), projected LiDAR points (as keys), and LiDAR points (as values), thereby capturing more accurate correspondences between LiDAR and camera data points. The model fundamentally learns and generates a correction matrix ($H^{\Delta}$) to refine the initial calibration, resulting in an improved matrix ($H^{*}$) for better LiDAR–camera alignment.
3.3 Common Feature Discriminator

### v4

@@ II Related Works
Extrinsic calibration methods for LiDAR–camera systems can generally be divided into two categories: target-based and targetless approaches. Target-based calibration relies on specially designed calibration targets and is thus commonly associated with manual, offline procedures. Moreover, these methods typically cannot handle real-time decalibrations, which are common in practical scenarios. In contrast, targetless methods extract features directly from natural scenes, making them well-suited for automatic, online calibration. These methods can be broadly categorized into motion-based, edge alignment-based, mutual information-based, and deep learning–based approaches.
II-1 Motion-based Calibration
Motion-based calibration methods leverage sensor movements or relative poses derived from visual and LiDAR odometry to compute extrinsic parameters [39, 26, 10]. For instance, Petek et al. [39] utilize odometry paths from each sensor, aligning them through non-linear optimization and dense 2D–3D matching. Park et al. [26] similarly derive closed-form calibration from relative sensor transformations. Despite effectiveness under certain conditions, these methods heavily rely on accurate odometry or SLAM estimations, which are susceptible to noise and degenerate in scenarios with limited sensor motion (e.g., minimal rotation). Hand-eye calibration–based methods [28] further exacerbate this issue, requiring multiple precise sensor poses, thus limiting their practical applicability, particularly for static installations.
II-2 Edge Alignment-based Calibration
Edge-based approaches attempt to align edges detected from LiDAR point clouds and camera images [5, 29, 30]. For example, Zhang et al. [5] transform the calibration problem into a cylindrical projection-based 2D–2D alignment task, while Li et al. [29] employ advanced edge extraction techniques such as the Segment Anything Model (SAM) combined with multi-frame filtering. However, reliably matching edges between sensors is inherently challenging due to modality differences—LiDAR captures sparse geometric structures, whereas camera edges reflect dense texture and lighting variations. These differences often lead to inaccurate feature correspondence and suboptimal calibration outcomes.
II-3 Mutual Information-based Calibration
Mutual information (MI)-based calibration methods utilize statistical relationships between LiDAR reflectance and camera image intensities [31, 40]. Pandey et al. [31], for instance, maximize mutual information between LiDAR reflectance intensities and camera images to find optimal alignment. Despite their conceptual elegance, these methods struggle due to significant variability in LiDAR reflectance caused by surface material differences and camera pixel intensity fluctuations arising from lighting changes, leading to inconsistent and less reliable calibration results.
II-4 Deep Learning-based Calibration
Deep learning-based methods have introduced neural networks to regress extrinsic calibration parameters directly [23, 1, 17, 25, 24]. Seminal works such as RegNet [23] and LCCNet [1] regress calibration offsets using projected LiDAR depth maps derived from initial calibration estimates. Similarly, Xiao et al. [17] utilize transformer-based architectures to refine feature correspondences. Zhu et al. [24] propose CalibDepth, which uses depth maps as a unified representation across modalities, integrating monocular depth estimation and sequence modeling to improve online calibration performance. However, these methods depend heavily on an initial calibration to project LiDAR point clouds into the image plane, forming a ”projected LiDAR depth map” that aligns sparse LiDAR data with dense image pixels. While this enables cross-modal feature correlation, it significantly limits generalizability—since the initial calibration is often manually provided or empirically estimated. Moreover, due to the sparsity of LiDAR data, the resulting depth maps are dominated by image features, effectively sidelining useful LiDAR-specific information and heavily biases calibration toward image modality. In addition, direct regression tasks pose significant computational challenges due to their unconstrained nature, further limiting real-time applicability. Alternative semantic segmentation-based methods [41] also face computational inefficiencies without substantial accuracy benefits over simpler object detection-based methods.
In summary, existing automatic, targetless, and online calibration methods commonly exhibit limitations including reliance on accurate odometry or sensor movement, difficulty in cross-modal edge matching, sensitivity to reflectance and illumination conditions, and computational inefficiencies. Furthermore, most existing methods fail to fully exploit the advances in object detection and feature extraction developed for LiDAR and camera data processing. To overcome these limitations, we propose CalibRefine, a fully automatic, targetless, and online calibration framework directly processing raw LiDAR point clouds and camera images, eliminating initial calibrations or elaborate preprocessing. Our approach integrates proven object detection algorithms and introduces a novel Common Feature Discriminator for robust cross-sensor correspondence matching. Furthermore, we employ a coarse-to-fine strategy combining iterative optimization and attention-driven refinement, enabling accurate and robust real-time calibration. By directly matching corresponding points across modalities, our approach facilitates a straightforward, one-shot, and end-to-end calibration process between the LiDAR and camera, significantly enhancing adaptability to real-world scenarios.
III Proposed Method
III-A Problem Formulation
Extrinsic calibration between sensors aims to unify detections from two different sensors into the same frame of reference or coordinate system, enabling the fusion of their respective detection information. LiDAR–camera extrinsic calibration is typically accomplished by solving for a transformation matrix that associates a point in the image pixel coordinate system (PCS) with its corresponding point in the LiDAR coordinate system (LCS). Since points in the LCS are 3D, while those in the PCS are 2D, most existing calibration methods rely on 3D-to-2D perspective projection. However, this approach has notable drawbacks. First, it requires the camera’s intrinsic matrix, adding the burden of intrinsic camera calibration, which is often performed manually, thus hindering fully automatic extrinsic calibration. Second, estimating the 3D-to-2D transformation matrix is computationally more complex and prone to instability. More importantly, for most practical applications, 3D-to-2D perspective calibration is unnecessary for achieving effective LiDAR–camera data fusion. A simpler 2D-to-2D projective calibration, where the 2D LiDAR plane is obtained by removing the Z-axis, is sufficient.
@@ III-A Problem Formulation
This simplification is justified for several reasons. The primary goal of calibration is to enable data fusion between the two sensors, such as associating 3D LiDAR point cloud clusters with image pixel regions for the same object. Achieving this does not require projecting the 3D LiDAR point cloud onto the image plane using a 3D-to-2D calibration matrix. Most existing methods adopt the 3D-to-2D approach as it draws from camera calibration practices that focus on 3D reconstruction. However, LiDAR–camera calibration is fundamentally different, as its focus is on data fusion, not reconstruction. By using 2D-to-2D projective calibration, where 2D LiDAR points are mapped to the image plane, corresponding 3D LiDAR points can still be retrieved without requiring a 3D-to-2D perspective transformation. Additionally, when projecting 3D LiDAR points onto the image plane, the LiDAR data effectively becomes 2D, resulting in the loss of LiDAR’s inherent 3D detection capabilities. Therefore, 3D-to-2D calibration does not offer additional benefits over 2D-to-2D projective calibration. Notably, while many existing methods emphasize projecting 3D LiDAR points onto the image plane, this should only serve as a visualization tool to intuitively present calibration performance, not as the calibration objective itself. The true goal of calibration should be the seamless and accurate fusion of sensor data.
Thus, we propose using planar projective transformation to achieve 2D-to-2D calibration between the 2D LiDAR plane and the camera image plane, as illustrated in Fig. 5. A planar projective transformation, or Homography, is an invertible linear transformation represented by a non-singular matrix $\mathbf{H}\in\mathbb{R}^{3\times 3}$ [42]. This transformation allows us to project a point in the LCS directly onto the camera image plane without requiring the camera intrinsic matrix. The relationship is expressed as:
@@
where $\hat{P_{l}}=(u_{l},v_{l})$ is the projection of a point $P_{l}=(x,y)$ in the LCS onto the camera image plane PCS. Notably, the objects or points on the 2D LiDAR plane and those on the camera image plane are derived from objects or points lying on a common plane (e.g., the ground plane), as shown in Fig. 5. This alignment justifies the use of 2D Homography for LiDAR–camera calibration, as it can be considered a planar homography induced by the common plane [43]. To solve for the Homography matrix, a set of $N$ points in the LCS and their corresponding points in the PCS is required. Although 4 points are theoretically sufficient, using more points allows optimization of the solution via a cost function that minimizes the geometric reprojection error [44, 42]. This error, which measures the alignment between $N$ projected LiDAR points ($\hat{P_{l}}=(\hat{u},\hat{v})$) and their corresponding 2D image pixel points ($P_{p}=(u,v)$), can be quantified either as the Average Euclidean Distance (AED)
@@
$\begin{split}\mathcal{E}_{\text{AED}}&=\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}\\ &=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{v}^{i})^{2}},\end{split}$ (2)
or as the Root Mean Square Error (RMSE)
@@
$\begin{split}\mathcal{E}_{\text{RMSE}}&=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}^{2}}\\ &=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left[(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{u}^{i})^{2}\right]}.\end{split}$ (3)
III-B Method Overview
We aim to develop a framework for LiDAR–camera online automatic targetless calibration that reduces human intervention, streamlines sensor integration, and ensures high precision in LiDAR–camera fusion applications. Our method comprises the following stages, as shown in Fig. 1: In Stage 1, established LiDAR and camera detectors are used to extract objects from each sensor’s data, capturing their center positions (bounding box centers for the camera detections and cluster centers for the LiDAR point cloud). These objects serve as training samples for a Common Feature Discriminator, which determines whether an image object and a LiDAR object correspond to the same entity by learning and comparing three distinct features (Relative Positions, Appearance Embeddings, and Classification Information) before concatenating them and passing them through a feed-forward neural network (FFN) classifier. In Stage 2, a homography transformation is applied to generate a coarse initial calibration matrix $H$ based on the identified object pairs, establishing a preliminary correspondence between the LiDAR and camera detections. Stage 3 refines this initial matrix iteratively by projecting LiDAR data onto the camera plane and creating additional point correspondences according to distance criteria, leading to a more precise calibration. Finally, in Stage 4, attention-based refinement employs a Vision Transformer to extract global distortion features from images, mitigating issues arising from non-planar surfaces, depth variations, and the absence of intrinsic matrix-based rectification. A cross-attention network computes weighted interactions among image pixels (queries), projected LiDAR points (keys), and LiDAR points (values), capturing more accurate correspondences between the two data sources. The model then learns a correction matrix $H^{\Delta}$ to refine the initial calibration, yielding an improved matrix $H^{*}$ for superior LiDAR–camera alignment.
III-C Common Feature Discriminator